# Skill-Enhanced Reinforcement Learning Acceleration from Demonstrations

### v1

@@ 4.2 Downstream Policy Training
In the downstream policy training stage, we aim to exploit the skill knowledge learned from the demonstration data, encoded by the skill prior network $q_{\psi}(\cdot)$ and decoder network $p_{\nu}(\cdot|z_{t})$, to accelerate the online RL process. To this end, we train a skill-based policy network $\pi_{\theta}(z_{t}|s_{t})$ for the downstream online RL task with skill-level behavior cloning.
Specifically, following the work [17], when interacting with the environment, we sample a skill $z_{t}$ from the skill policy network $\pi_{\theta}(\cdot|s_{t})$ given the current state $s_{t}$. The skill $z_{t}$ is then decoded to an action sequence $a_{t:t+H-1}$ using the skill decoder $p_{\nu}(\cdot|z_{t})$ to guide the RL agent to reach state $s_{t+H}$ in $H$ steps. The cumulative reward over the $H$ steps, i.e., the $H$-step reward, can be collected from the environment as $\tilde{r}_{t}=\sum_{t}^{t+H}r_{t}$. With the online skill-based transition data $D=\{(s_{t},z_{t},\tilde{r}_{t},s_{t+H})\}$, we deploy a Skill-based soft actor-critic (SSAC) algorithm utilized in [17] to conduct skill-based policy learning with skill-level behavior cloning. SSAC extends soft actor-critic (SAC) [37] to learn the skill policy function network $\pi_{\theta}(z_{t}|s_{t})$ with the support of a skill-based soft Q-function network $Q_{\phi}(s_{t},z_{t})$. It utilizes a KL-divergence regularization term to enforce the skill policy function $\pi_{\theta}(z_{t}|s_{t})$ to clone the behavior of the pre-trained skill prior network $q_{\psi}(z_{t}|s_{t})$.
@@
Input: Skill prior network $q_{\psi}(z_{t}|s_{t})$, and decoder $p_{\nu}(a_{t:t+H-1}|z_{t})$,  $\eta\in(0,1)$
Initialize: Replay buffer $D$, skill policy $\pi_{\theta}(z_{t}|s_{t})$, and skill-based $Q$ value functions
Output: Trained skill policy network $\pi_{\theta}(z_{t}|s_{t})$
@@
6:         Update buffer: $D\leftarrow D\cup\{s_{t},z_{t},\tilde{r}_{t},s_{t+H}\}$
@@
11:         Update skill-policy function and Q-function
12:     end for
13:end for
In particular, SSAC learns the skill policy function network by maximizing the following regularized expected skill-based Q-value:
@@ 4.3 Skill-Level Data Enhancement
Collecting expert demonstrations can be challenging and expensive, due to the involvement of human experts [8]. The scarcity of the limited expert data however hinders the robust and effective learning of skills. To alleviate the problem, in addition to incorporating general random demonstration data during the prior training stage, we further propose a Skill-level Data Enhancement (SDE) technique to augment the skill-level data in both the skill prior training stage and the downstream policy training stage, improving the robustness of learning at the skill-level.
Conventional data augmentation is typically applied to the input data, e.g., state observations. It is not applicable to the action space since actions (continuous or discrete) cannot be easily re-represented or rescaled. By using a latent variable model to learn skills as latent representations for behaviours captured by action sequences, our proposed approach provides a new augmentation space at the skill level, without interfering with the real action space. Specifically, we propose to augment our skill level data with Gaussian noise altered versions as follows. For each skill embedding $z_{t}$, we can add a Gaussian noise altered version $\hat{z}_{t}=z_{t}+\eta\epsilon$ into learning, where $\epsilon\sim\mathcal{N}^{m}(0,1)$ is a Gaussian noise vector sampled from a $m$-dimensional independent standard Gaussian distribution, $m$ is the dimension of the skill embedding, and $\eta\in(0,1)$ is a very small scaling factor. We then enforce $z_{t}$ and $\hat{z}_{t}$ correspond to the same action sequence, aiming to achieve stable and differentiable representations for different behaviors in the skill embedding space, and enhance the robustness of skill learning.
@@
$\hat{L}_{rec}(\nu,\mu)=\alpha\,\mathbb{E}_{{\bf a}_{t}\sim A_{\pi_{e}}}\ell_{ls}\left(\hat{\bf a}_{t}\sim p_{\nu}\big{(}\!\cdot|\hat{z}_{t}\sim\eta\epsilon\!+\!q_{\mu}(\cdot|{\bf a}_{t})\big{)},{\bf a}_{t}\right)$ (10)
@@
In the downstream skill policy training stage, SDE is realized efficiently by augmenting the skill-based transition data. For each observed skill-based transition $\{s_{t},{z}_{t},\tilde{r}_{t},s_{t+H}\}$, we produce an altered skill vector and add an augmenting transition $\{s_{t},\hat{z}_{t},\tilde{r}_{t},s_{t+H}\}$ to buffer $D$, without any extra interaction with the environment. The goal is to make the skill-policy network more robust to small variations in the skill embedding space, accelerating the learning process. The online skill policy training process augmented with SDE is outlined in Algorithm 2.
@@ 5.1 Experimental Setting
We conduct experiments with four demonstration-guided tasks with long-horizon and sparse rewards in four different environments that are commonly used for skill learning: Maze, Kitchen, Mis-aligned Kitchen from the D4RL datasets [38] and CALVIN [18] from CALVIN challenge [39]. Maze is a navigation environment, in which a point mass agent is required to find a path between a randomly initialized starting point and a goal point [28]. Kitchen is a robotic manipulation environment in which a robotic arm completes a sequence of four sub-tasks (Microwave-Kettle-Bottom Burner-Light) [11]. Mis-aligned Kitchen is modified from the Kitchen environment with a different task sequence (Microwave-Light-Slide Cabinet-Hinge Cabinet) [28, 18]. CALVIN is a Language-Conditioned Policy Learning challenge[39], which has been adapted for skill learning in SkiMo [18], requiring the RL agent to complete a sequence of four sub-tasks in order: Open Drawer, Turn on Lightbulb, Move Slider Left, and Turn on LED.
@@
We compared the proposed SeRLA with two state-of-the-art skill-based methods, SPiRL [17] and SkiMo [18], which train skill priors for the downstream tasks.
Figure 1: The comparison results on the four long-horizon sparse-reward tasks (Maze, Kitchen, Mis-aligned Kitchen, and CALVIN) are presented in the figure. Each plot presents the average per-trajectory return (i.e., reward) v.s. environment steps for each environment during the downstream training process.
@@ 5.2 Experimental Results
We conducted experiments on the four environments to compare the proposed full method, SeRLA, and its variant without SDE, SeRLA-w/o-SDE, with the other two skill-based methods, SPiRL and SkiMo. The skills were learned on the demonstration data prior to the downstream task and the reward was evaluated in the downstream RL learning process over $10^{7}$ environment steps. The maximum trajectory reward for Maze environment is $1$, while for Kitchen, Mis-aligned Kitchen, and CALVIN environments, it is $4$. The experimental results are presented in Figure 1. The four plots on the left side report results (return v.s. environment steps) on the four environments separately. We can see that the proposed SeRLA-w/o-SDE largely outperforms the baseline SPiRL across all the four environments, especially on Maze, Mis-aligned Kitchen and CALVIN, which validates the effectiveness and contribution of the proposed PU Learning component and SSAC algorithm in extracting and deploying skill knowledge. The proposed full approach SeRLA further boosts the performance over SeRLA-w/o-SDE with notable gains across the four plots, which verifies the effectiveness of the proposed skill-level data enhancement technique. SeRLA also outperforms the model-based state-of-the-art SkiMo and produces best results on Kitchen and Mis-aligned Kitchen. On the other two environments, Maze and CALVIN, SeRLA yields comparable overall performance to SkiMo throughout the downstream RL training, while producing the best results in the early training stage.
@@ 6 Conclusion
In this paper, we proposed a novel two-stage skill-level learning method SeRLA to exploit offline demonstration data and accelerate downstream RL tasks. SeRLA deploys skill-level adversarial PU learning to learn reusable skills from both limited expert demonstration data and large low-cost demonstration data. Then a skill-based soft actor-critic algorithm is deployed to utilize the learned skill prior knowledge and accelerate the online downstream RL through skill-based behavior cloning. The proposed approach conveniently provides a new augmentation space at the skill level without interfering with the real action space, which enables novel skill-level data enhancement (SDE) in both training stages. Our experimental results on four benchmark environments demonstrate that SeRLA outperforms two state-of-the-art skill learning methods, SPiRL and SkiMo, especially in the early downstream training stage. Given the promise and effectiveness of our simple but innovative skill-level data enhancement technique, we plan to further investigate various data augmentation strategies in the high level latent skill space in the future.

### v2

@@ 4.2 Downstream Policy Training
In the downstream policy training stage, we aim to exploit the skill knowledge learned from the heterogeneous demonstration data, encoded by the skill prior network $q_{\psi}(\cdot)$ and decoder network $p_{\nu}(\cdot|z_{t})$, to accelerate the online RL process. To this end, we train a skill-based policy network $\pi_{\theta}(z_{t}|s_{t})$ for the downstream online RL task with skill-level behavior cloning.
Specifically, when interacting with the environment, we sample a skill $z_{t}$ from the skill policy network $\pi_{\theta}(\cdot|s_{t})$ given the current state $s_{t}$. The skill $z_{t}$ is then decoded to an action sequence $a_{t:t+H-1}$ using the skill decoder $p_{\nu}(\cdot|z_{t})$ to guide the RL agent to reach state $s_{t+H}$ in $H$ steps. The cumulative reward over the $H$ steps, i.e., the $H$-step reward, can be collected from the environment as $\tilde{r}_{t}=\sum_{t}^{t+H}r_{t}$. With the online skill-based transition data $D=\{(s_{t},z_{t},\tilde{r}_{t},s_{t+H})\}$, we deploy a Skill-based soft actor-critic (SSAC) algorithm [31] to conduct skill-based policy learning with skill-level behavior cloning. SSAC extends soft actor-critic (SAC) [16] to learn the skill policy function network $\pi_{\theta}(z_{t}|s_{t})$ (i.e., actor) with the support of a skill-based soft Q-function network $Q_{\phi}(s_{t},z_{t})$ (i.e., critic). In particular, SSAC learns the skill policy function network by maximizing the following regularized expected skill-based Q-value:
Different from the standard SAC algorithm, which regularizes the policy function through its KL-divergence from a uniform distribution, the KL-divergence regularization term in SSAC enforces the skill policy function $\pi_{\theta}(z_{t}|s_{t})$ to clone the behavior of the pre-trained skill prior network $q_{\psi}(z_{t}|s_{t})$. From the value function’s perspective, the KL-regularizer in Eq.(9) can be interpreted as imposing a penalty on the skill-level Q-value in the context of behavior cloning. The soft skill-level Q-value function, $Q_{\phi}(s_{t},z_{t})$, is trained to minimize the following soft Bellman residual objective:
$J_{Q}(\phi)=\!\!\!\!\!\!\!\mathop{\mathbb{E}}_{(\begin{subarray}{c}s_{t},z_{t},\tilde{r}_{t},\\ s_{t+H}\end{subarray})\sim D}\left[\frac{1}{2}\big{(}Q_{\phi}(s_{t},z_{t})-(\tilde{r}_{t}\!+\!\gamma V_{\bar{\phi}}(s_{t+H}))\big{)}^{2}\right],$ (10)
which is computed on the data collected from the online interaction with the environment, and $V$ denotes the state value function. To facilitate the learning of $Q_{\phi}$, we keep a stabilized Q-function $Q_{\bar{\phi}}(\cdot,\cdot)$ by computing its parameters $\bar{\phi}$ from $\phi$ using an exponential moving average (EMA): $\bar{\phi}\leftarrow\tau\phi+(1-\tau)\bar{\phi}$, where $\tau\in(0,1]$ is a momentum parameter. Based on [25], we then compute $V_{\bar{\phi}}(s_{t+H})$ from the penalized and stabilized soft Q-function $Q_{\bar{\phi}}(s_{t+H},z_{t+H})$ as:
$V_{\bar{\phi}}(s_{t+H})$ $=\mathbb{E}_{z_{t+H}\sim\pi_{\theta}}\big{[}Q_{\bar{\phi}}(s_{t+H},z_{t+H})$
$-\kappa\mathcal{L}_{KL}(\pi_{\theta}(z_{t+H}|s_{t+H}),q_{\psi}(z_{t+H}|s_{t+H}))\big{]},$ (11)
which incorporates the behavior cloning regularization into the Q-function learning as well, efficiently utilizing the prior knowledge acquired from the demonstration data. The main process of the SSAC algorithm for downstream policy training is outlined in Algorithm 2.
@@
Input: Skill prior network $q_{\psi}(\cdot)$, decoder $p_{\nu}(\cdot)$, $\eta\in(0,1)$
Initialize: Replay buffer $D$, skill policy $\pi_{\theta}(\cdot)$, critics $Q_{\phi}$ and $Q_{\bar{\phi}}$
Output: Trained skill policy network $\pi_{\theta}(\cdot)$
@@
6:   Update buffer: $D\leftarrow D\cup\{s_{t},z_{t},\tilde{r}_{t},s_{t+H}\}$% SDE augmentation steps in line 7-8:
@@
11:   Update policy parameters $\theta$ by maximizing Eq.(9)
12:   Update Q-function parameters $\phi$ by minimizing Eq.(10)
13:   EMA update $\bar{\phi}\leftarrow\tau\phi+(1-\tau)\bar{\phi}$
14:  end for
15:end for
@@ 4.3 Skill-Level Data Enhancement
Collecting expert demonstrations can be challenging and expensive, due to the involvement of human experts [5]. The scarcity of the limited expert data however hinders the robust and effective learning of skills. To alleviate the problem, in addition to incorporating general, low-cost demonstration data during the prior training stage, we further propose a Skill-level Data Enhancement (SDE) technique to augment the skill-level data in both the skill prior training stage and the downstream policy training stage, improving the robustness of learning at the skill-level.
Conventional data augmentation is typically applied to the input data, e.g., state observations. It is not applicable to the action space since actions (continuous or discrete) cannot be easily re-represented or rescaled. By using a latent variable model to learn skills as latent representations for behaviours captured by action sequences, our proposed approach provides a new augmentation space at the skill level, without interfering with the real action space. Specifically, we propose to augment our skill level data with Gaussian noise altered versions as follows. For each skill embedding $z_{t}$, we can add a Gaussian noise altered version $\hat{z}_{t}=z_{t}+\eta\epsilon$ into the learning process, where $\epsilon\sim\mathcal{N}^{m}(0,1)$ is a Gaussian noise vector sampled from a $m$-dimensional independent standard Gaussian distribution, $m$ is the dimension of the skill embedding, and $\eta\in(0,1)$ is a very small scaling factor. We then enforce $z_{t}$ and $\hat{z}_{t}$ correspond to the same action sequence, aiming to achieve stable representations for different behaviors in the skill embedding space, and enhance the robustness of skill learning.
@@
$\!\!\!\hat{L}_{rec}(\nu,\mu)=\alpha\!\!\!\!\!\!\mathop{\mathbb{E}}_{{\bf a}_{t}\sim A_{\pi_{e}}}\!\!\!\ell_{ls}\!\left(\hat{\bf a}_{t}\!\sim\!p_{\nu}\big{(}\cdot|\hat{z}_{t}\sim\eta\epsilon\!+\!q_{\mu}(\cdot|{\bf a}_{t})\big{)},{\bf a}_{t}\right),$ (12)
@@
In the downstream skill policy training stage, SDE is realized efficiently by augmenting the skill-based transition data. For each observed skill-based transition $\{s_{t},{z}_{t},\tilde{r}_{t},s_{t+H}\}$, we produce an altered skill vector $\hat{z}_{t}$ from $z_{t}$ and add an augmenting transition $\{s_{t},\hat{z}_{t},\tilde{r}_{t},s_{t+H}\}$ to buffer $D$, without any extra interaction with the environment. The goal is to make the skill-policy network more robust to small variations in the skill embedding space, accelerating the learning process. The online skill policy training process augmented with SDE is outlined in Algorithm 2.
@@ 5 Experiment
Figure 2: The four environments used in the experiments. Top row: Maze and Kitchen. Bottow row: Mis-aligned Kitchen and CALVIN. In Maze, a point agent navigates from a starting point (green) to the target point (red). In each of the other three robotic manipulation environments, a robot arm completes four different sub-tasks in order.
Figure 3: The comparison results on the four long-horizon sparse-reward tasks (Maze, Kitchen, Mis-aligned Kitchen, and CALVIN) are presented in the figure. Each plot presents the average per-trajectory return (i.e., reward) v.s. environment steps for each environment during the downstream training process. The results were collected through 5 random seeds.
Figure 4: The plot summarizes the results across the four environments and presents the mean normalized return averaged over the four environments v.s. environment steps. The results were collected through 5 random seeds.
@@ 5.1 Experimental Setting
We conduct experiments with four demonstration-guided tasks with long-horizon and sparse rewards in four different environments that are commonly used for skill learning: Maze, Kitchen, Mis-aligned Kitchen, and CALVIN [36], which are shown in Figure 2. The first three environments (Maze, Kitchen, and Mis-aligned Kitchen) are adapted from the D4RL datasets [12], while the last environment (CALVIN) is adapted from the CALVIN challenge [28]. Maze is a navigation environment, in which a point mass agent is required to find a path between a fixed starting point and a goal point. It is modified by randomly initializing the starting point of the agent around the original starting point [32]. The RL agent receives a sparse reward of 1 only when it reaches the close neighbor of the goal point and receives a reward of 0 otherwise. Kitchen is a robotic manipulation environment, in which a robotic arm completes a sequence of four sub-tasks (Microwave–Kettle–Bottom Burner–Light) [15]. The RL agent receives a sparse binary reward only when it completes each sub-task in sequence, and can obtain four reward scores if it completes all four sub-tasks in order. Mis-aligned Kitchen is modified from the Kitchen environment with a different task sequence (Microwave–Light–Slide Cabinet–Hinge Cabinet) [32, 36]. Unlike in the original Kitchen setting, here the subtask order in expert demonstrations is misaligned with that in the downstream task, which makes skill learning more challenging. CALVIN is a robotic manipulation environment designed for Language-Conditioned Policy Learning [28]. The environment has been adapted for skill learning in SkiMo [36], requiring the RL agent to complete a sequence of four sub-tasks in order: Open Drawer, Turn on Lightbulb, Move Slider Left, and Turn on LED. The agent receives a binary reward signal for each sub-task it completes in order, with the maximum value of per-trajectory reward as 4.
Table 1: This table shows the percentage improvements achieved by variants of SeRLA with Skill-level Data Enhancement (SDE) over the ablation baseline SeRLA-w/o-SDE. The results are the average percentage increases of each method over SeRLA-w/o-SDE across the whole downstream training stage.
  Average Increase Maze Kitchen Mis-aligned Kitchen CALVIN
SeRLA_SDE (skill) $22\pm 0.30 0.0$$41\pm 0.010 0.0$$032\pm 0.025$ $0.35\pm 0.30$
SeRLA_SDE (downstream) $15\pm 0.19 0.0$$24\pm 0.052 0.0$$021\pm 0.018$ $0.29\pm 0.26$
SeRLA (full) $26\pm 0.34 0.0$$87\pm 0.066 0.0$$038\pm 0.026$ $0.41\pm 0.35$
@@
We compared the proposed SeRLA with two state-of-the-art skill-based methods, which train skill priors for the downstream tasks:
  • SPiRL [31] learns a skill prior using a deep latent variable model from expert demonstration data to guide policy training in the downstream task.
  • SkiMo [36] is a model-based method. It learns reusable skills and a skill dynamic model in offline training, and selects the optimal skill in the downstream training using long-term model-based planning.
Implementation Details
We built SeRLA on top of SPiRL [31] for skill prior training, and used the official implementations of the two comparison methods, SPiRL [31] and SkiMo [36]. The skill horizon was fixed to $H=10$ and the skill embedding dimension is set as $m=10$. In skill-level adversarial PU learning, the discriminator $\mathcal{D}_{\zeta}$ is implemented as an MLP that has two 256-unit hidden layers with ReLU activations, followed by a Xavier-initialized linear output head. We set the positive class prior to $\lambda=0.5$, the relaxation slack variable to $\xi=0$, and the PU loss trade-off parameter to $\rho=0.1$. The trade-off parameter $\beta$ for the Gaussian prior regularizer is set to the same value as in SPiRL. For the SDE, the scaling factor is fixed at $\eta=0.01$, and the trade-off parameter for the augmenting loss is set as $\alpha=0.1$. Downstream policy learning follows the SAC formulation [16] with $\tau=0.005$ and $\gamma=0.99$, while $\kappa$ is treated as a dual variable and updated during training.
Datasets
The expert dataset $D_{\pi_{e}}$ contains demonstrations obtained either from human experts or from fully trained RL agents. We used the version aggregated in the SkiMo [36] repository, which contains expert demonstrations collected in the four environments: Maze, Kitchen, Mis-aligned Kitchen, and CALVIN. In Maze, the expert demonstration data is originally collected in the work of SPiRL [32], consisting of 3,046 trajectories. In Kitchen and Mis-aligned Kitchen, the expert data is originally from the D4RL dataset [12], comprising 603 trajectories. In CALVIN, the expert data is from the CALVIN challenge [28], comprising 1,239 trajectories. In each environment, the general, low-cost demonstration data $D_{\pi}$ was collected using an RL agent that is pre-trained with $10^{5}$ timesteps of interaction with the environment, starting from scratch. In comparison to the full training regime of $10^{7}$ timesteps, the pre-trained RL agent is significantly undertrained, operating with a near-random policy. However, it has learned some basic behaviors that support exploration of the environments with minimal training cost. In each environment, the general low-cost demonstration data contains ten times as many trajectories as the expert data.
@@ 5.2 Experimental Results
We conducted experiments on the four environments (Maze, Kitchen, Mis-aligned Kitchen, and CALVIN) to compare the proposed full method, SeRLA, and its variant without SDE, SeRLA-w/o-SDE, with the other two skill-based comparison methods, SPiRL and SkiMo. The skills were learned on the heterogeneous demonstration data prior to the downstream task and the reward was evaluated in the downstream RL learning process over $10^{6}$ environment steps. The maximum trajectory reward for the Maze environment is $1$, while for the Kitchen, Mis-aligned Kitchen, and CALVIN environments, it is $4$. The experimental results are presented in Figure 3 and Figure 4. In Figure 3, the four plots report results (return v.s. environment steps) on the four environments separately. We can see that the proposed SeRLA-w/o-SDE largely outperforms the baseline SPiRL across all the four environments, especially on Maze, Mis-aligned Kitchen and CALVIN, which validates the effectiveness and contribution of the proposed PU Learning component in extracting useful skill knowledge from heterogeneous demonstration data. The proposed full approach SeRLA further improves performance over SeRLA-w/o-SDE, yielding notable gains across all the four plots, which verifies the effectiveness of the proposed skill-level data enhancement technique. SeRLA also outperforms the model-based state-of-the-art SkiMo and produces best results on Kitchen and Mis-aligned Kitchen. On the other two environments, Maze and CALVIN, SeRLA yields comparable overall performance to SkiMo throughout the downstream RL training, while producing the best results in the early training stage.
To present a more illustrative overall comparison across the four environments, we evaluate the performance of each method by calculating its mean normalized return across all four environments, with reward from each environment being normalized to the range of $[0,1]$ based on its maximum possible reward [8]. The mean normalized return is obtained by taking the average normalized reward across the four environments, and plotted in Figure 4. The results clearly show that SeRLA outperforms SkiMo in the early training stage, achieving the best overall performance among all the comparison methods. This validates the effectiveness of the proposed approach.
5.3 Ablation Study
The previous comparison results have shown that the full approach SeRLA outperforms SeRLA-w/o-SDE and validated the contribution of the SDE technique. We further conducted an ablation study to assess the impact of the SDE technique at the two training stages separately. We experimented with two variants of SeRLA: (1) SeRLA_SDE (skill), which denotes SeRLA with SDE applied only on the skill prior training; (2) SeRLA_SDE (downstream), which denotes SeRLA with SDE applied only on the downstream policy training. We evaluate these two variants and the full approach against the baseline SeRLA-w/o-SDE that entirely drops SDE from both training stages, and record the percentage increase of their performance values over that of SeRLA-w/o-SDE.
We collected the average increase of each method over the entire downstream training stage and reported the results in Table 1. The results show that adding SDE to either training stage separately can produce performance gains within the proposed SeRLA framework. In comparison, SDE is much more effective when applied on the skill prior training stage, particularly on the Maze and CALVIN environments. This validates the efficacy of SDE in boosting the robustness of skill embedding learning, which consequently improves the performance of downstream RL tasks. When applying SDE on both training stages, there are still some marginal performance increases than applying it on each stage separately. These results again validate the contribution of SDE to the proposed approach.
@@ 6 Conclusion
In this paper, we proposed a novel two-stage skill-level learning method SeRLA to exploit heterogeneous offline demonstration data and accelerate downstream RL tasks. SeRLA deploys skill-level adversarial PU learning to learn reusable skills from both limited expert demonstration data and large low-cost demonstration data. Then a skill-based soft actor-critic algorithm is deployed to utilize the learned skill prior knowledge and accelerate the online downstream RL through skill-based behavior cloning. The proposed approach conveniently provides a new augmentation space at the skill level without interfering with the real action space, which enables novel skill-level data enhancement (SDE) in both training stages. Our experimental results on four benchmark environments demonstrate that SeRLA outperforms two state-of-the-art skill learning methods, SPiRL and SkiMo, especially in the early downstream training stage.